OpenStudy (anonymous):

1 cg is equal to medal and fan! 10^3 hg. 10^3 g. 10^-2 mg. 10^-2 g.

jimthompson5910 (jim_thompson5910):

hint: 1 cg = 1 centigram 'centi' means 'hundredth' (think of a penny, one 'cent', which is one hundredth of a dollar)

OpenStudy (anonymous):

So either C or D , im not sure how to figure out if its mg or g? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

1 cg = 1 centigram 1 cg = 1 hundredth of a gram (not 1 hundredth of a mg)

jimthompson5910 (jim_thompson5910):

centigram = centi+gram (not centi+mg)

OpenStudy (anonymous):

so D?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

since 10^(-2) = 1/100 = 0.01

jimthompson5910 (jim_thompson5910):

np